How many liters of 15% acid and 33% acid should be mixed to make 40 liters of 21% acid solution ?

Answer by Alan3354(69443)

How many liters of 15% acid and 33% acid should be mixed to make 40 liters of 21% acid solution ?
==============================
Here's an example.
-----------------
If an order is written for 500mls of 15% solution and all you have in stock is 10% and 70%. How many mls of each will you need to compound order?
-----------------
Do these like this:
t = 10%
s = 70%
-------
t + s = 500 (total solution)
10t + 70s = 500*15 (total ingredient)